# Animate Anyone: Consistent and Controllable Image-to-Video Synthesis for Character Animation

### v2

@@ Abstract
Character Animation aims to generating character videos from still images through driving signals. Currently, diffusion models have become the mainstream in visual generation research, owing to their robust generative capabilities. However, challenges persist in the realm of image-to-video, especially in character animation, where temporally maintaining consistency with detailed information from character remains a formidable problem. In this paper, we leverage the power of diffusion models and propose a novel framework tailored for character animation. To preserve consistency of intricate appearance features from reference image, we design ReferenceNet to merge detail features via spatial attention. To ensure controllability and continuity, we introduce an efficient pose guider to direct character’s movements and employ an effective temporal modeling approach to ensure smooth inter-frame transitions between video frames. By expanding the training data, our approach can animate arbitrary characters, yielding superior results in character animation compared to other image-to-video methods. Furthermore, we evaluate our method on benchmarks for fashion video and human dance synthesis, achieving state-of-the-art results.
@@ 1 Introduction
Our model is trained on an internal dataset of 5K character video clips. Fig. 1 shows the animation results for various characters. Compared to previous methods, our approach presents several notable advantages. Firstly, it effectively maintains the spatial and temporal consistency of character appearance in videos. Secondly, it produces high-definition videos without issues such as temporal jitter or flickering. Thirdly, it is capable of animating any character image into a video, unconstrained by specific domains. We evaluate our method on two specific human video synthesis benchmarks (UBC fashion video dataset[55] and TikTok dataset[18]), using only the corresponding training datasets for each benchmark in the experiments. Our approach achieves state-of-the-art results. we also compare our method with general image-to-video approaches trained on large-scale data and our approach demonstrates superior capabilities in character animation. We envision that Animate Anyone could serve as a foundational solution for character video creation, inspiring the development of more innovative and creative applications.
@@ 2.3 Diffusion Model for Human Image Animation
Image animation[35, 37, 31, 60, 36, 57, 54, 6] aims to generate images or videos based on one or more input images. In recent research, the superior generation quality and stable controllability offered by diffusion models have led to their integration into human image animation. PIDM[3] proposes texture diffusion blocks to inject the desired texture patterns into denoising for human pose transfer. LFDM[26] synthesizes an optical flow sequence in the latent space, warping the input image based on given conditions. LEO[47] represents motion as a sequence of flow maps and employs diffusion model to synthesize sequences of motion codes. DreamPose[19] utilizes pretrained Stable Diffusion model and proposes an adapter to model the CLIP and VAE image embeddings. DisCo[45] draws inspiration from ControlNet, decoupling the control of pose and background. Despite the incorporation of diffusion models to enhance generation quality, these methods still grapple with issues such as texture inconsistency and temporal instability in their results. Moreover, there is no method to investigate and demonstrate a more generalized capability in character animation.
@@ 4.1 Implementations
To demonstrate the applicability of our approach in animating various characters, we collect 5K character video clips (2-10 seconds long) from the internet to train our model. We employ DWPose[52] to extract the pose sequence of characters in the video, including the body and hands, rendering it as pose skeleton images following OpenPose[5]. Experiments are conducted on 4 NVIDIA A100 GPUs. In the first training stage, individual video frames are sampled, resized, and center-cropped to a resolution of $768{\times}768$. Training is conducted for 30,000 steps with a batch size of 64. In the second training stage, we train the temporal layer for 10,000 steps with 24-frame video sequences and a batch size of 4. Both learning rates are set to 1e-5. During inference, we rescale the length of the driving pose skeleton to approximate the length of the character’s skeleton in the reference image and use a DDIM sampler for 20 denoising steps. We adopt the temporal aggregation method in [41], connecting results from different batches to generate long videos. For fair comparison with other image animation methods, we also train our model on two specific benchmarks (UBC fashion video dataset[55] and TikTok dataset[18]) without using additional data, as will be discussed in Section 4.3.
Figure 3: Qualitative Results. Given a reference image (the leftmost image), our approach demonstrates the ability to animate diverse characters, encompassing full-body human figures, half-length portraits, cartoon characters, and humanoid figures. The illustration showcases results with clear, consistent details, and continuous motion. More video results are available in supplementary material.
@@ 4.2 Qualitative Results
Fig. 3 demonstrates that our method can animate arbitrary characters, including full-body human figures, half-length portraits, cartoon characters, and humanoid characters. Our approach is capable of generating high-definition and realistic character details. It maintains temporal consistency with the reference images even under substantial motion and exhibits temporal continuity between frames. More video results are available in the supplementary material.
@@ 4.3 Comparisons
To demonstrate the effectiveness of our approach compared to other image animation methods, we evaluate its performance in two specific benchmarks: fashion video synthesis and human dance generation. For quantitative assessment of image-level quality, SSIM[48], PSNR[16] and LPIPS[58] are employed. Video-level evaluation uses FVD[42] metrics.
Figure 4: Qualitative comparison for fashion video synthesis. DreamPose and BDMM exhibit shortcomings in preserving fine-textured details of clothing, whereas our method excels in maintaining exceptional detail features. More comparison could be found in supplementary material.
@@
Fashion Video Synthesis. Fashion Video Synthesis aims to turn fashion photographs into realistic, animated videos using a driving pose sequence. Experiments are conducted on the UBC fashion video dataset, which consists of 500 training and 100 testing videos, each containing roughly 350 frames. The quantitative comparison is shown in Tab. 1. Our result outperforms other methods, particularly exhibiting a significant lead in video metric. Qualitative comparison is shown in Fig. 4. For fair comparison, we obtain results of DreamPose without sample finetuning using its open-source code. In the domain of fashion videos, there is a stringent requirement for fine-grained clothing details. However, the videos generated by DreamPose and BDMM fail to maintain the consistency of clothing details and exhibit noticeable errors in terms of color and fine structural elements. In contrast, our method produces results that effectively preserves the consistency of clothing details.
Figure 5: Qualitative comparison between DisCo and our method. DisCo’s generated results display problems such as pose control errors, color inaccuracy, and inconsistent details. In contrast, our method demonstrates significant improvements in addressing these issues.
@@
Human Dance Generation. Human Dance Generation focuses on animating images in real-world dance scenarios. We utilize the TikTok dataset, comprising 340 training and 100 testing single human dancing videos (10-15 seconds long). Following DisCo’s dataset partitioning and utilizing an identical test set (10 TikTok-style videos), we conduct a quantitative comparison presented in Tab. 2, and our method achieves the best results. For enhanced generalization, DisCo incorporates human attribute pre-training, utilizing a large number of image pairs for model pre-training. In contrast, our training is exclusively conducted on the TikTok dataset, yielding results superior to DisCo. We present qualitative comparison with DisCo in Fig. 5. Given the scene’s complexity, DisCo’s pipeline involves an extra step with SAM[23] to generate the human foreground mask. Conversely, our approach showcases that, even without explicit learning of human mask, the model can grasp the foreground-background relationship from the subject’s motion, negating the necessity for prior human segmentation. Additionally, during intricate dance sequences, our model stands out in maintaining visual continuity throughout the motion and exhibits enhanced robustness in handling diverse character appearances.
Figure 6: Qualitative comparison with image-to-video methods. Current image-to-video methods struggle to generate substantial character movements and face challenges in maintaining long-term consistency with character image features.
@@ 4.4 Ablation study
To demonstrate the effectiveness of the ReferenceNet design, we explore alternative designs, including (1) using only the CLIP image encoder to represent reference image features without integrating ReferenceNet, (2) initially finetuning SD and subsequently training ControlNet with the reference image. (3) integrating the above two designs. Experiments are conducted on the UBC fashion video dataset. As shown in Fig. 7, visualizations illustrate that ReferenceNet outperforms the other three designs. Solely relying on CLIP features as reference image features can preserve image similarity but fails to fully transfer details. ControlNet does not enhance results as its features lack spatial correspondence, rendering it inapplicable. Quantitative results are also presented in Tab. 3, demonstrating the superiority of our design.
Figure 7: Ablation study of different design. Only our ReferenceNet ensures the consistent preservation of details in the character’s appearance.
@@
Table 3: Quantitative comparison for ablation study.
5 Limitations
The limitations of our method can be summarized in three aspects: First, similar to many visual generation models, our model may struggle to generate highly stable results for hand movements, sometimes leading to distortions and motion blur. Second, since images provide information from only one perspective, generating unseen parts during character movement is an ill-posed problem which encounters potential instability. Third, due to the utilization of DDPM, our model exhibits a lower operational efficiency compared to non-diffusion-model-based methods.
6 Conclusion
In this paper, we present Animate Anyone, a character animation framework capable of transforming character photographs into animated videos controlled by a desired pose sequence while ensuring consistent appearance and temporal stability. We propose ReferenceNet, which genuinely preserves intricate character appearances and we also achieve efficient pose controllability and temporal continuity. Our approach not only applies to general character animation but also outperforms existing methods in specific benchmarks. Animate Anyone serves as a foundational method, with the potential for future extension into various image-to-video applications.

### v3

@@ Abstract
Character Animation aims to generating character videos from still images through driving signals. Currently, diffusion models have become the mainstream in visual generation research, owing to their robust generative capabilities. However, challenges persist in the realm of image-to-video, especially in character animation, where temporally maintaining consistency with detailed information from character remains a formidable problem. In this paper, we leverage the power of diffusion models and propose a novel framework tailored for character animation. To preserve consistency of intricate appearance features from reference image, we design ReferenceNet to merge detail features via spatial attention. To ensure controllability and continuity, we introduce an efficient pose guider to direct character’s movements and employ an effective temporal modeling approach to ensure smooth inter-frame transitions between video frames. By expanding the training data, our approach can animate arbitrary characters, yielding superior results in character animation compared to other image-to-video methods. Furthermore, we evaluate our method on image animation benchmarks, achieving state-of-the-art results.
@@ 1 Introduction
Our model is trained on an internal dataset of 5K character video clips. Fig. 1 shows the animation results for various characters. Compared to previous methods, our approach presents several notable advantages. Firstly, it effectively maintains the spatial and temporal consistency of character appearance in videos. Secondly, it produces high-definition videos without issues such as temporal jitter or flickering. Thirdly, it is capable of animating any character image into a video, unconstrained by specific domains. We evaluate our method on three specific human video synthesis benchmarks (UBC fashion video dataset[59], TikTok dataset[20] and Ted-talk dataset[39]), using only the corresponding training datasets for each benchmark in the experiments. Our approach achieves state-of-the-art results. We also compare our method with general image-to-video approaches trained on large-scale data and our approach demonstrates superior capabilities in character animation. We envision that Animate Anyone could serve as a foundational solution for character video creation, inspiring the development of more innovative and creative applications.
@@ 2.3 Diffusion Model for Human Image Animation
Image animation[37, 39, 33, 64, 38, 61, 57, 7, 25, 13], aims to generate images or videos based on one or more input images. In recent research, the superior generation quality and stable controllability offered by diffusion models have led to their integration into human image animation. PIDM[3] proposes texture diffusion blocks to inject the desired texture patterns into denoising for human pose transfer. LFDM[28] synthesizes an optical flow sequence in the latent space, warping the input image based on given conditions. LEO[49] represents motion as a sequence of flow maps and employs diffusion model to synthesize sequences of motion codes. DreamPose[21] utilizes pretrained Stable Diffusion model and proposes an adapter to model the CLIP and VAE image embeddings. DisCo[47] draws inspiration from ControlNet, decoupling the control of pose and background. Despite the incorporation of diffusion models to enhance generation quality, these methods still grapple with issues such as texture inconsistency and temporal instability in their results. Moreover, there is no method to investigate and demonstrate a more generalized capability in character animation.
@@ 3.3 Training Strategy
Figure 3: Qualitative Results. Given a reference image (the leftmost image of each group), our approach demonstrates the ability to animate diverse characters, encompassing full-body human figures, half-length portraits, cartoon characters, and humanoid figures. The illustration showcases results with clear, consistent details, and continuous motion.
@@ 4.1 Implementations
To demonstrate the applicability of our approach in animating various characters, we collect 5K character video clips from the internet to train our model. We employ DWPose[55] to extract pose sequence of characters in the video, including body and hands, rendering it as pose skeleton images following OpenPose[6]. Experiments are conducted on 4 NVIDIA A100 GPUs. In the first training stage, individual video frames are sampled, resized, and center-cropped to a resolution of $768{\times}768$. Training is conducted for 30,000 steps with a batch size of 64. In the second training stage, we train the temporal layer for 10,000 steps with 24-frame video sequences and a batch size of 4. Both learning rates are set to 1e-5. During inference, we rescale the length of the driving pose skeleton to approximate the length of the character’s skeleton in the reference image and use a DDIM sampler for 20 denoising steps. We adopt the temporal aggregation method in [43], connecting results from different batches to generate long videos. For fair comparison with other methods, we also train our model on three specific benchmarks (UBC fashion video dataset[59], TikTok dataset[20] and Ted-talk dataset[39]) without using additional data, as will be discussed in Section 4.3.
@@ 4.2 Qualitative Results
Fig. 3 demonstrates that our method can animate arbitrary characters, including full-body human figures, half-length portraits, cartoon characters, and humanoid characters. Our approach is capable of generating high-definition and realistic character details. It maintains temporal consistency with the reference images even under substantial motion and exhibits temporal continuity between frames.
@@ 4.3 Comparisons
To demonstrate the superiority of our approach, we evaluate its performance in three specific benchmarks: fashion video synthesis, human dance generation and talking gesture generation. We also conduct a baseline that combines Stable Diffusion, ControlNet, IP-Adapter[56] and AnimateDiff, named SD-I2V. For quantitative assessment of image-level quality, SSIM[50], PSNR[18] and LPIPS[62] are employed. Video-level evaluation uses FVD[44] metrics.
Figure 4: Qualitative comparison for fashion video synthesis. Other methods exhibit shortcomings in preserving fine-textured details of clothing, whereas our method excels in maintaining exceptional detail features.
@@
Fashion Video Synthesis. Experiments are conducted on the UBC fashion video dataset. The quantitative comparison is shown in Tab. 1. Our result outperforms other methods, particularly exhibiting a significant lead in video metric. Qualitative comparison is shown in Fig. 4. For fair comparison, we obtain results of DreamPose without sample finetuning using its open-source code. In the domain of fashion videos, there is a stringent requirement for fine-grained clothing details. However, other methods fail to maintain the consistency of clothing details and exhibit noticeable errors in terms of color and fine structural elements. In contrast, our method produces results that effectively preserve the consistency of clothing details.
Figure 5: Qualitative comparison between DisCo and our method. DisCo displays problems such as pose control errors, color inaccuracy, and inconsistent details. In contrast, our method demonstrates significant improvements in addressing these issues.
@@
Human Dance Generation. We conduct experiments on the TikTok dataset. We conduct a quantitative comparison presented in Tab. 2, and our method achieves the best results. For enhanced generalization, DisCo incorporates human attribute pre-training, utilizing a large number of image pairs for model pre-training. In contrast, our training is exclusively conducted on the TikTok dataset, yielding results superior to DisCo. We present qualitative comparison with DisCo in Fig. 5. During intricate dance sequences, our model stands out in maintaining visual continuity throughout the motion and exhibits enhanced robustness in handling diverse character appearances.
Figure 6: Qualitative comparison on Ted-talk dataset. Our model is capable of generating more accurate and clear results.
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
MRAA[39] 0.826 33.86 0.160 82.8
TPSMM[64] 0.830 33.81 0.157 80.7
Disco[47] 0.754 31.25 0.193 223.5
SD-I2V 0.773 32.11 0.179 158.3
Ours 0.832 33.91 0.159 80.5
Table 3: Quantitative comparison on Ted-talk dataset.
Talking Gesture Generation. We also evaluate our method on Ted-talk dataset. Results are shown in Fig 6 and Tab 3. Our approach significantly outperforms DisCo and SD-I2V. MRAA and TPSMM employ GT images as driving signals (video reconstruction), while we achieve better results using only pose information. On other two evaluated benchmarks (UBC with more intricate clothing textures and TikTok with more complex human movements), the performance of MRAA and TPSMM falls far behind our method.
Figure 7: Qualitative comparison with image-to-video methods, which struggle to generate substantial character movements and face challenges in maintaining long-term appearance consistency.
@@ 4.4 Ablation study
Image Condition Modeling. To demonstrate the effectiveness of our image condition modeling, we explore alternative designs, including 1) using only the CLIP image encoder to represent reference image features without integrating ReferenceNet, 2) initially finetuning SD and subsequently training ControlNet with the reference image. 3) integrating the above two designs. Experiments are conducted on UBC fashion video dataset. As shown in Fig. 8, visualizations illustrate that ReferenceNet outperforms the other three designs. Solely relying on CLIP features as reference image features can preserve image similarity but fails to fully transfer details. ControlNet does not enhance results as its features lack spatial correspondence, rendering it inapplicable. Quantitative results are also presented in Tab. 4, demonstrating the superiority of our design.
Figure 8: Ablation study of different design. ReferenceNet ensures consistent preservation of details in character’s appearance.
Details of ReferenceNet Design. To demonstrate the effectiveness of our ReferenceNet design, we conduct experiments: 1) Replacing UNet (SD weights) with ResNet (ImageNet weights). 2) Replacing spatial-attention with feature-concatenation. Quantitative result is shown in Tab 5. Our design achieves the optimal performance. Experiment 1) shows the necessity of utilizing SD weights. There exists a certain gap between ImageNet’s image features and the implicit features in SD. Utilizing features from SD enhances the integration of conditioning information within the same feature space during the generation process. Experiment 2) demonstrates the necessity of spatial-attention which enables the denoising UNet to effectively integrate detailed image features from ReferenceNet.
@@
Table 4: Ablation study for image condition modeling.
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
ImageNet weights 0.901 36.21 0.084 165.4
Feature-concat 0.909 36.53 0.071 132.8
Ours 0.931 38.49 0.044 81.6
Table 5: Ablation study for ReferenceNet design.
Temporal Modeling. We conduct two experiments to assess the effectiveness of proposed temporal modeling method: 1) do not apply temporal layer, directly concatenating images temporally to create a video. 2) do not apply the two-stage training, directly training the entire network. Quantitative results are presented in Tab. 6. The absence of temporal layer results in noticeable texture sticking and inter-frame jitter, leading to a significant decrease in FVD metric. When the two-stage training is not employed, metrics related to image quality experience a decline. We attribute this to the fact that, when optimizing over multiple frames simultaneously, the network tends to focus more on the overall temporal visual coherence, thereby weakening attention to the details of each individual frame. The adoption of the two-stage training method ensures both the quality of generated video frames and temporal smoothness.
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
w/o Temporal Layer 0.925 38.28 0.049 176.7
w/o Two Stage Training 0.917 38.01 0.056 89.3
Ours 0.931 38.49 0.044 81.6
Table 6: Ablation study of temporal modeling.
5 Discussion and Conclusion
Limitations. Our model may struggle to generate stable results for hand movements, sometimes leading to distortions and motion blur. Besides, since images provide information from only one perspective, generating unseen parts during character movement is an ill-posed problem which encounters potential instability. Third, due to the utilization of DDPM, our model exhibits a lower operational efficiency compared to non-diffusion-model-based methods.
Potential Impact. The proposed method may be used to produce fake videos of individuals, which can be detected using some face anti-spoofing techniques[5, 51, 58].
Conclusion. In this paper, we present Animate Anyone, a framework capable of transforming character photographs into animated videos controlled by a desired pose sequence. We propose ReferenceNet, which genuinely preserves intricate character appearances and we also achieve efficient pose controllability and temporal continuity. Our approach not only applies to general character animation but also outperforms existing methods.